OpenStudy (i_love_my_nieces):

Find the next three terms of the sequence. 3, 6, 9, 12, . . . 15, 18, 20 15, 19, 21 15, 18, 21 14, 18, 21

OpenStudy (anonymous):

It's only going up by 3, so 3+3 = 6 + 3 = 9 + 3 = 12 +3 =...

OpenStudy (anonymous):

Do you get it?

OpenStudy (anonymous):

Just add three to the previous number.

OpenStudy (i_love_my_nieces):

C?

OpenStudy (anonymous):

Yes :)

OpenStudy (i_love_my_nieces):

Thank you!